OpenStudy (abhisar):

The relation between time t and distance x is \(\sf t = ax^² + bx\) where a and b are constants. The acceleration is

OpenStudy (abhisar):

I don't know how to do implicit differentiation.....

OpenStudy (solomonzelman):

This is kind of an implicit expression. \(\large\color{black}{ \displaystyle t=ax^2+bx }\) I would first try to remodel it as a distance a function of time. \(\large\color{black}{ \displaystyle 0=ax^2+bx-t }\) \(\large\color{black}{ \displaystyle x=\frac{-b\pm\sqrt{b^2+4at}}{2a} }\) where x is distance and t is time.

OpenStudy (solomonzelman):

So this is kind of a \(\large\color{black}{ \displaystyle x(t)=\frac{-b\pm\sqrt{b^2+4at}}{2a} }\) where x(t) is a distance at a particular time t.

OpenStudy (solomonzelman):

Acceleration is a 2nd derivative (with respect to t)

OpenStudy (abhisar):

Yep I know that but i was not able to differentiate it correctly, would you mind checking my work?

OpenStudy (solomonzelman):

Yes, sure....

OpenStudy (solomonzelman):

oh, lets hold a sec and re-write the function better....

OpenStudy (abhisar):

Ok, So i was heading like this, differentiating the equation on both side by dt, \(\sf \frac{d}{dt}t= ax.\frac{dx}{dt}+b\frac{dx}{dt}\\ 1=axv + bv\)

OpenStudy (solomonzelman):

and for even less confusion \(\large\color{black}{ \displaystyle x(t)=\frac{-b+\sqrt{b^2+4at}}{2a} }\) \(\large\color{black}{ \displaystyle x(t)=\frac{-b}{2a}+\frac{\sqrt{b^2+4at}}{2a} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x(t)=\frac{-b}{2a}+\frac{1}{2a}\cdot\sqrt{b^2+4at} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x'(t)=0+\frac{1}{2a}\cdot(b^2+4at)^{1/2~\color{red}{-1}} \color{red}{\times (1/2)}}\)

OpenStudy (solomonzelman):

kind of like that, if it makes sense to you. (we can do a u-sub, if it is too messy for you)

OpenStudy (solomonzelman):

oh, i frgot ththe chain

OpenStudy (abhisar):

Yep I am getting it.

OpenStudy (abhisar):

@SolomonZelman but can you check what i did and confirm if it is correct?

OpenStudy (solomonzelman):

oh, sure.. let me look back..

OpenStudy (abhisar):

Ok, So i was heading like this, differentiating the equation on both side by dt, \(\sf \frac{d}{dt}t= ax.\frac{dx}{dt}+b\frac{dx}{dt}\\ 1=axv + bv\)

OpenStudy (solomonzelman):

yah that seems better than my way

OpenStudy (solomonzelman):

((and your v is your velocity)) now, a second deriv. is going to be accelearation.

OpenStudy (abhisar):

Yes, exactly but I am finding a different story in one of my book for this question. Lemme copy that for you.

OpenStudy (abhisar):

Differentiate the equation on both sides wrt dt which will give \(\sf 1 = 2axv + bv\)

OpenStudy (solomonzelman):

wrt?

OpenStudy (solomonzelman):

with respect?

OpenStudy (abhisar):

wrt = with respect to time

OpenStudy (solomonzelman):

yeah that is what you should do, you differentiate with respect to time (or with respect to t, and your function is now velocity v, so you do dv/dt)

OpenStudy (abhisar):

Ok but I did the same and got 1 = axv + bv

OpenStudy (solomonzelman):

wait, you are saying that you got this for the 2nd deriv. or what ?

OpenStudy (abhisar):

For first derivative I got 1 = axv + bv but the book says 1=2axv + bv. How?

OpenStudy (solomonzelman):

oh, the power of 2 by x²

OpenStudy (solomonzelman):

t=ax²+bx (d/dt) t= (d/dt) ax+ (d/dt) bx (dt/dt)= 2 (dx/dt) ax²⁻¹ + (dx/dt) b

OpenStudy (solomonzelman):

you are differentiating after all when you mutliply times d/dt... so this is what you get, you would get a coefficient of 3 and a power of 2, if you have ax³, and you would get a coefficient of n, and power of n-1, if you had axⁿ

OpenStudy (solomonzelman):

This is not just •C on every term in the equation. Playing with derivatives gets a little different.....

OpenStudy (abhisar):

ok, so if i want to differentiate say \(\sf x^2\) with respect to time I get d/dt \(\sf x^2\) = 2x/dt = 2v ?

OpenStudy (solomonzelman):

2x and times the derivative of x', if you are differentiating with respect to x. In this case this x' was the velocity (and namely, the dx/dt, or as you named it - which is typical - "v")

OpenStudy (solomonzelman):

This is going to the implicit differentiation, that when you find dy/dx of a function in a form of f(x,y)=.... where the function is not defined explicity (has y's and x's mixed up - roughly speaking, (and that is why, as well) doesn't pass vertical line test)

OpenStudy (solomonzelman):

Want an implicit differention example, real quick?

OpenStudy (abhisar):

Yes, I think that's my problem. I don't know how to such problems.

OpenStudy (solomonzelman):

Lets do some implicit differentiation... it will be easy if you know basic properties of regular (not implicit differentiation).

OpenStudy (abhisar):

Yes I know basics.

OpenStudy (solomonzelman):

yeah, so i will do example or 2-3....

OpenStudy (abhisar):

ok

OpenStudy (solomonzelman):

say i got: 4x+2y²x=4y (y is output, x is input) I differentiate all parts applying my rules, but... dy/dx (4x¹) = dy/dx (4x)=4x¹⁻¹ = 4x⁰ = 4 dy/dx (4y) = 4 • y' (same as by 4x, but we multiply times y', since y is a function of x.) Where does this come from? I will tell you. So say you wanted to take a derivative of a huge function f(x), you will have a chain rule - multiply times f'(x) (correct?) Here y' plays the same role. You are taking the derivative of this huge function f(x) (and y is here to denote a function f(x)), so just like by taking a derivative of a function f(x) you get a chain rule of f'(x), so is when you differentiate [a function] y, you get a chain rule of y'. dy/dx (2y²x) = x • dy/dx (2y²) + 2y² • dy/dx (x) (taking the product rule, and will apply the same chain rule principal by every derivative of y) it becomes: dy/dx (2y²x) = x • (2• 2y²⁻¹) • y' + 2y² • 1 see the y' here that i explained before? (and derivative of x, with respect to x, is just 1) simplifying: dy/dx (2y²x) = x(4y)y' + 2y² So overall, when we differentiate the: 4x+2y²x=4y (finding dy/dx) we are going to get: 4 + [ x(4y)y' + 2y² ] = 4 y'

OpenStudy (solomonzelman):

sorry for long reply. (there is one step in implicit differentiation to solve for y' (just algebraic manipulations)

OpenStudy (solomonzelman):

the important part is the chain rule of y' note. (take you time)

OpenStudy (abhisar):

ok, so what is y'?

OpenStudy (solomonzelman):

y' is the chain rule. y differens from x, in that y is [denoting] a function of x. here are some equivalent statements: dy/dx[ f(x) ] = f'(x) \(\Rightarrow\) dy/dx[y]=y' dy/dx[ ( f(x) )³ ] = 3f(x)² • f'(x) \(\Rightarrow\) dy/dx[y³]=3y²•y'

OpenStudy (solomonzelman):

So basically y' is f'(x)

OpenStudy (abhisar):

Ok, I think I got it. So Here I need to apply chain rule. For example if i want to differentiate the original equation, I will do it like this, \(\sf \frac{d}{dt}(t) = \frac{dx}{dt}.(2ax)+ b.\frac{dx}{dt}\)

OpenStudy (abhisar):

\(\sf \Rightarrow 1 = 2axV + bV)\)

OpenStudy (solomonzelman):

yes

OpenStudy (abhisar):

Now, 1/v = 2ax + b Differentiating it one more time \(\sf -v^{-2}.a = 2a.V + 0 \\ \Rightarrow a = -2av^3\)

OpenStudy (solomonzelman):

why do you have .a on the left side? and with respect to what are you differentiating?

OpenStudy (abhisar):

With time

OpenStudy (abhisar):

dv/dt = a (acceleration)

OpenStudy (abhisar):

The constant a on the right hand side is different..

OpenStudy (abhisar):

Hehe, I think that's correct.

OpenStudy (solomonzelman):

lets see..... What i was gonna do from the beginning is: \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle x(t)=\frac{-b+\sqrt{b^2+4at}}{2a} } }\) \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle x(t)=\frac{-b}{2a}+\frac{\sqrt{b^2+4at}}{2a} } }\) \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle x'(t)=0+\frac{4a}{4a\sqrt{b^2+4at}} } }\) \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle x'(t)=\frac{1}{\sqrt{b^2+4at}} } }\) \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle x''(t)=\left(\sqrt{b^2+4at}\right)^{-1/2~~\color{red}{-1}} ~\cdot (1/2)4a} }\) \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle x''(t)=2a\left(\sqrt{b^2+4at}\right)^{-3/2}} }\) so, \(\large\color{black}{ \displaystyle \large\color{black}{ \displaystyle A=\frac{2a}{\sqrt{(b^2+4at)^3}}} }\)

OpenStudy (solomonzelman):

This is how I was going to do it, without any use of creativity or brain. Simply solve for x in terms of t, and find the derivative twice (both times doing dx/dt)

OpenStudy (solomonzelman):

A is the acceleration

OpenStudy (abhisar):

Ummm.. I have four options 2av^2 2av^3 2abv^3 2b^2v^3

OpenStudy (abhisar):

Which means i need to find the value in terms of v (velocity)

OpenStudy (solomonzelman):

yeah...

OpenStudy (solomonzelman):

1=2axv+bv 1/v=2ax+b differentiating ... (dv/dt)1/v=(dx/dt)2ax+(d/dt)b -v\(^{-2}\) • \(\color{red}{\rm A}\) =2a•x' \(\color{red}{\rm A}\) =-2v²a•x' this x' is v. \(\color{red}{\rm A}\) =-2v²a•v \(\color{red}{\rm A}\) =-2v³a

OpenStudy (solomonzelman):

but I got negative same thing...

OpenStudy (abhisar):

Yes that's correct. In option it's positive because they are asking Retarding force.

OpenStudy (abhisar):

Thanks for bearing with me c:

OpenStudy (solomonzelman):

Kind of an unusual for me. I would tend to find acceleration with respect to time (to know how particle accelerates at paricular time).... got it thorugh, tho'. next time won';t play basketball before going on here... hehe.. you are welcome, you were great!

OpenStudy (abhisar):

Lel, thank you for time and help, you were life saver today ;)

OpenStudy (solomonzelman):

Okay... good luck